If f(1) = 3 and f(n) = -2f(n-1) + 1, then what is f(5)?
Answer by greenestamps(13200)



This is a recursively defined sequence, where each number is found from the preceding number based on a pattern. The recursive definition

f(n) = -2f(n-1) + 1

means each term (f(n), term "n") is found by multiplying the previous term (f(n-1), the preceding term) by -2 and adding 1.

Given: f(1)=3

f(2)=-2*f(1)+1 = -2(3)+1 = -6+1 = -5
f(3)=-2*f(2)+1 = -2(-5)+1 = 10+1 = 11

Continue using the pattern to find f(5).